OpenStudy (anonymous):

The given measurements may or may not determine a triangle. If not, then state that no triangle is formed. If a triangle is formed, then use the Law of Sines to solve the triangle, if it is possible, or state that the Law of Sines cannot be used. C = 38°, a = 19, c = 10

OpenStudy (anonymous):



OpenStudy (anonymous):

This triangle cannot exist because by the Law of Sines:\[\frac{ \sin 38 }{ 10 } = \frac{ \sin A }{ 19 }\]but that would give a sin A that is greater than "1" and that is impossible.

OpenStudy (anonymous):

sin A = (19)(sin 38)/(10) = 1.17 Can't have that.

OpenStudy (anonymous):

All good now @coolkat4 ?

OpenStudy (anonymous):

can it not be solved with law od sines or no triangle could be formed period?

OpenStudy (anonymous):

Both because the law of sine would hold for all triangles. It cannot be solved by the law of sines AND the triangle cannot exist.

OpenStudy (anonymous):

That 38 degree angle is just too big

OpenStudy (anonymous):

okay thank you!